I need help in figuring out how to set this problem up and solving same:
Five divided by the sum of a number and 7, minus the quotient of 3 and the difference of the number and 7 is equal to 6 times the reciprocal of the difference of the number squared and 49. What is the number?
Thank You To Any and All ...

Answer by edjones(8007)

5/(x+7) - 3/(x-7) = 6/(x^2-49)
5(x-7)-3(x+7)=6 Multiply each side by (x+7)(x-7)
5x-35-3x-21=6
2x=62
x=31
.
Ed